OpenStudy (anonymous):

Clear fractions or decimals, solve, and check . please show work for checking. 1/2 + 4m =3m – 5/2

OpenStudy (anonymous):

m=-3

OpenStudy (anonymous):

I got the answer to the problem but i need help when checking!

OpenStudy (anonymous):

1/2+4(-3)=(-3)-5/2 1/2-12=-9-5/2

OpenStudy (anonymous):

replace m by -3 and get \[\frac{1}{2}+4\times -3=3\times -3+\frac{5}{2}\]

OpenStudy (anonymous):

to check your answer just plug in -3 for m and see if lhs = rhs

OpenStudy (anonymous):

\[\frac{1}{2}-12=-9-\frac{5}{2}\] \[-11\tfrac{1}{2}=-9-2\tfrac{1}{2}\] \[-11\tfrac{1}{2}=-11\tfrac{1}{2}\]

OpenStudy (anonymous):

shouldn't it equal 0=0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

left hand side should equal right hand side , thats all

OpenStudy (anonymous):

If you start adding 5/2 to both sides, you are following the same logic you used when you worked the problem. This makes it less likely to catch any mistakes with your checking. You need to figure out what 1/2 - 12 is, and also what 9 - 5/2 is. Work it all the way until you have either something like 1 = 1

OpenStudy (anonymous):

sorry - i dont follow that

OpenStudy (anonymous):

okay thanks immyrep!

OpenStudy (anonymous):

thats ok